### v3

# Generalists vs. Specialists: Evaluating Large Language Models for Urdu

Abstract
In this paper, we compare general-purpose models, GPT-4-Turbo and Llama-3-8b, with special-purpose models—XLM-Roberta-large, mT5-large, and Llama-3-8b—that have been fine-tuned on specific tasks. We focus on seven classification and seven generation tasks to evaluate the performance of these models on Urdu language. Urdu has 70 million native speakers, yet it remains underrepresented in Natural Language Processing (NLP). Despite the frequent advancements in Large Language Models (LLMs), their performance in low-resource languages, including Urdu, still needs to be explored. We also conduct a human evaluation for the generation tasks and compare the results with the evaluations performed by GPT-4-Turbo, Llama-3-8b and Claude 3.5 Sonnet. We find that special-purpose models consistently outperform general-purpose models across various tasks. We also find that the evaluation done by GPT-4-Turbo for generation tasks aligns more closely with human evaluation compared to the evaluation the evaluation done by Llama-3-8b. This paper contributes to the NLP community by providing insights into the effectiveness of general and specific-purpose LLMs for low-resource languages.
Generalists vs. Specialists: Evaluating Large Language Models for Urdu
Samee Arif  and Abdul Hameed Azeemi  and Agha Ali Raza {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk Lahore University of Management Sciences
Awais Athar awais@ebi.ac.uk EMBL European Bioinformatics Institute
1 Introduction
In recent years the introduction of LLMs including GPT (Brown et al., 2020; OpenAI, 2024) and Llama (Touvron et al., 2023a, b) has led to a significant advancement in NLP. However, expanding the reach of NLP to low-resource languages is crucial for advancing multilingual AI systems and promoting technological inclusivity. Urdu, with over 70 million native speakers, stands as a significant yet underserved language in the NLP domain (Blasi et al., 2022).
For this study, we classify LLMs into two distinct categories:
  1. 1.
    Generalists: General-purpose models capable of performing a wide variety of tasks. We will use GPT-4-Turbo (abbreviated as GPT) trained on dataset up to Dec 2023 and Llama-3-8b (abbreviated as Llama) as the generalist models.
  2. 2.
    Specialists: Special-purpose models fine-tuned to perform specific tasks. We use XLM-Roberta-large (abbreviated as XLM-R) (Conneau et al., 2020), mT5-large (abbreviated as mT5) (Xue et al., 2021), and a fine-tuned version of Llama-3-8b (abbreviated as Llama-FT) as the specialist models.
We present a comprehensive evaluation of generalist and specialist models for classification and generation tasks, exploring their strengths and limitations. Table 1 outlines the sub-tasks associated with both categories, illustrating the scope of our evaluation.
Classification Generation
Sentiment Analysis Question Answering
Abuse Detection Summarization
Sarcasm Detection Paraphrasing
Fake News Detection Transliteration
Topic Classification Translation (en-ur)
Part-of-Speech Tagging Translation (ur-en)
Named-entity Recognition AI Assistant
Table 1: Sub-tasks for classification and generation.
In this paper, we aim to answer: (1) how each category of models performs on Urdu language tasks, and (2) which model type is more effective in practical applications for Urdu-speaking users. Specifically, we seek to determine if the added specialization of the specialist models translates into significant performance gains over the generalist models in the defined tasks.
Our contributions can be summarized as follows:
  1. 1.
    We fine-tune XLM-R, mT5, and Llama on classification and generation tasks to optimize their performance for the defined tasks.
  2. 2.
    We compare the performance of both generalist and specialist models on a smaller, controlled test set consisting of min(1000, len(dataset["test"])) samples, and provide a detailed comparison using various metrics. The exact size of each test set is given in Appendix D. Given the inherent challenges of working with low-resource languages like Urdu, our test set size reflects the current limitations in available data.
  3. 3.
    We conduct a human evaluation of the outputs from the generation tasks and compare these results with the automated evaluations performed by GPT, Llama and Claude 3.5 Sonnet (abbreviated as Claude).
The code, model outputs, and the human and LLM evaluations are publicly available on GitHub [^1] https://github.com/sameearif/Generalists-vs-Specialists.
2 Related Work
In recent years there has been a growing interest in the performance of LLMs across various languages. The MEGA benchmark (Ahuja et al., 2023) evaluates 16 NLP datasets across 70 languages. They compare the performance of BLOOMZ, GPT models, and State of the Art (SOTA) non-autoregressive models. MEGAVERSE Ahuja et al. (2024) builds on top of the MEGA benchmark and evaluates the non-English capabilities of GPT-3.5-Turbo, GPT-4, PaLM2, Gemini-Pro, Mistral, Llama-2, and Gemma. IndicGenBench (Singh et al., 2024) evaluates LLMs on user-facing generation tasks across a set of 29 Indic languages. They perform evaluation on both proprietary and open-source LLMs including GP-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM, and Llama. They cover the following tasks: cross-lingual summarization, machine translation, and cross-lingual question-answering. Zhao et al. (2024) conduct an evaluation of Llama’s response quality based on LLM-Eval (Zhang et al., 2023), a benchmark comprising instruction tasks from 17 categories. Mujadia et al. (2024) presents a comprehensive translation evaluation using Llama for English and Indian languages. They found that LLM-based evaluator achieves a comparable score with human judgement.
Khondaker et al. (2023a) presents a in-depth evaluaton of BLOOMZ, ChatGPT and specialist models AraT5 and MARBERTv2. They evaluated these models for Arabic on 44 distinct language understanding and generation tasks. They also present a comparison between the human evaluation and GPT-4 evaluation. Additionally, they observed that the English prompt works better than the Arabic prompt. Abdelali et al. (2024) provides a benchmark of LLMs against SOTA models for Arabic NLP. They evaluated GPT-3.5-Turbo, GPT-4, BLOOMZ, and Jais-13b-Chat across 33 tasks using 61 publicly available datasets, resulting in 330+ experimental setups. They observed that SOTA models generally outperform LLMs in zero-shot learning, larger models with few-shot learning techniques significantly reduce the performance gaps.
Existing multilingual NLP benchmarks often lack extensive language-specific evaluation and comparison against task-specific models. Tahir et al. (2024) address this gap by evaluating GPT-3.5-Turbo, Llama-2-7B-Chat, and Bloomz (3B and 7B) across 14 tasks using 15 Urdu datasets in a zero-shot setting, comparing their performance against task-specific models. Their findings reveal that task-specific models (Support Vector Machine (SVM), Decision Tree, and m-BERT, etc) generally outperform the mentioned LLMs in Urdu NLP tasks with zero-shot learning. However, they do not perform few-shot prompting, chain-of-thought prompting, or LLM fine-tuning.
3 Datasets
3.1 Classification
Sentiment Analysis. We use the Urdu IMDB sentiment analysis dataset [^2] https://github.com/urduhack/resources/releases/tag/imdb_urdu_reviews_v1.0.0, which is a translated version of the original IMDB dataset (Maas et al., 2011). It is translated using Google Translator, and comprises 50,000 movie reviews.
Abuse Detection. We use the dataset by Akhter et al. (2020) which has 2,171 entries and the dataset by Amjad et al. (2022) which consists of 3,502 entries.
Sarcasm Detection. For this task we use Urdu Sarcastic Tweets Dataset (Khan and Najeeb, 2023) which consists of 19,955 tagged tweets.
Fake News Detection. We use the fake news dataset by Amjad et al. (2020) which is a mixture of real and translated data. It comprises 1,300 labeled news articles.
Topic Classification. For this task, we use a dataset [^3] https://github.com/mwaseemrandhawa/Urdu-News-Headline-Dataset consisting of 137,161 news headlines categorized into the following topics: business, entertainment, health, politics, science, sports, world and other.
Part-of-Speech (PoS) Tagging. For this task, we use the Universal Dependencies (Nivre et al., 2020) dataset, which consists of 5,130 sentences, annotated with a PoS tag for every word. For GPT and Llama, we wrap the word for which we want to predict the PoS tag in <hl> tags. The structure of the data is given in Figure 1.
Figure 1: PoS Data Structure for Llama
Named Entity Recognition (NER). We use the dataset [^4] https://huggingface.co/datasets/mirfan899/urdu-ner available on Hugging Face which consists of 33,748 sentences, annotated with a NER tag for every word. For GPT and Llama, we applied the same data structuring method as done in the PoS tagging task.
3.2 Generation
Given our resource constraints, we could not set the maximum sequence length for model training to the longest entries in each dataset. Instead, we filter the datasets to include only those entries that fell within a manageable maximum length of 2048. This approach allowed us to optimize the training process and ensure efficient use of computational resources while still maintaining a representative sample of the data.
Question-Answering. We use three datasets for question-answering: (1) UQA (Arif et al., 2024a), consisting of 88,829 answerable questions; (2) UQuAD [^5] https://github.com/ahsanfarooqui/UQuAD---Urdu-Question-Answer-Dataset/tree/main, containing 139 questions; and (3) Wiki-UQA [^6] https://huggingface.co/datasets/uqa/Wiki-UQA, a manually generated dataset from Wikipedia articles, comprising 210 questions.
Summarization. For this task, we use the XSUMUrdu (Munaf et al., 2023) dataset, selecting a subset of 76,626 entries based on the maximum length used during model training.
Paraphrasing. For paraphrasing we use the dataset [^7] https://huggingface.co/datasets/mwz/ur_para available on Hugging Face. We select 387,004 entries from the dataset based on the maximum length used while training the models.
Transliteration. We use the Dakshina dataset (Roark et al., 2020) from which we select 11,464 sentences based on the maximum length used while training the models.
Translation. We use OPUS-100 (Zhang et al., 2020) (Tiedemann, 2012) for English-to-Urdu and Urdu-to-English translation. We select 755,526 sentences from this dataset based on the maximum length used while training the models.
AI Assistant. We use UrduAssistant (Khalil, 2023) dataset which is translated to Urdu from English and has 67,017 prompts.
4 Methodology
4.1 Experimental Design
We utilize a controlled test set consisting of min(1000, len(dataset["test"])) samples for each task except AI assistant, ensuring that the evaluation is both comprehensive and cost-efficient. For AI assistant we do human evaluation of 50 samples. We ensure that the test dataset is as balanced as possible, with an effort to achieve equal representation of different classes within each task. We use Macro-$F_{1}$ Score to evaluate all the classification tasks, SQuAD $F_{1}$ (Rajpurkar et al., 2018) Rajpurkar et al. (2016) to evaluate question-answering, SacreBLEU (Post, 2018) for paraphrasing, transliteration and translation, ROUGE-L (with word-level tokenization) Lin (2004) for summarization and Wins (the number of times a model is ranked one by the evaluator) for AI assistant evaluation.
We fine-tune Llama and XLM-R for each classification task separately, ensuring that each model is specifically optimized for its respective task. Similarly, for generation tasks, we fine-tune the mT5 and Llama separately for each task. For the mT5 models, we use a learning rate of $5e^{-5}$. The Llama models are fine-tuned with a learning rate of $2e^{-4}$ for both generation and classification tasks. In the case of XLM-R, we use a learning rate of $5e^{-6}$. We use LoRA (Hu et al., 2021) to fine-tune int4 quantized Llama. The batch size, number of epochs each model is trained for and LoRA config for Llama is given in Appendix E. After fine-tuning all the models we perform evaluation on the test dataset.
To evaluate the performance of the generalist models, we design a series of experiments. We use GPT and Llama as our generalist models. GPT is chosen due its top performance on the LMSYS chatbot arena (Chiang et al., 2024) as of March 1st, 2024 when we started our research. Our experimental setup is as follows: We conduct experiments under zero-shot, three-shot, and six-shot settings for both generation and classification tasks. The examples for the three-shot and six-shot scenarios are selected from the training dataset of each task. Specifically, the examples are carefully selected by a human expert to ensure a representative and balanced sample, avoiding any unintended bias in the selection process. We also conduct experiments with Chain-of-Thought (CoT) reasoning for classification tasks in the six-shot setting. We create CoT reasoning for the six selected examples from the training dataset. These examples along with their CoT is given as a few-shot prompt to the model. Figure 2 presents an example of generated CoT reasoning. For all the evaluations, the temperature and nucleus sampling for GPT-4 is set to 1.0, which is the default for the GPT API. For Llama, the temperature is set to 0.6 and nucleus sampling is set to 0.9, i.e., the values used in the original code-base for Llama [^8] https://github.com/meta-llama/llama3/blob/main/example_chat_completion.py.
Figure 2: CoT example from abuse detection
4.2 Prompt Design
Khondaker et al. (2023b) observe in their ChatGPT evaluation for Arabic that an English prompt performs better than the Arabic one. Therefore, following this study, we decide to use English prompts for our evaluations of Urdu tasks as well. In the prompt templates, the following placeholders are used:
  1. 1.
    ROLE: It specifies the persona of the LLM. For example, it could be sentiment classifier, abuse detector, sarcasm detector, etc.
  2. 2.
    TASK DESCRIPTION: It provides a brief description of what the task is and what the model is expected to do.
  3. 3.
    LABEL LIST: It lists the possible labels the model can assign to the input text. For example, it could be [’positive’, ’negative’] for sentiment analysis.
Figure 3 shows the prompt template for the classification task without CoT, and Figure 4 shows an example prompt for it. Figure 5 shows the CoT prompt template for the classification task. Figure 6 presents the prompt template for the generation task, and Figure 7 provides an example of it. Appendix G contains all the prompts for classification and generation tasks.
Figure 3: Classification Prompt Template
Figure 4: Classification Prompt Example
Figure 5: Classification Prompt Template (CoT)
Figure 6: Generation Prompt Template
Figure 7: Generation Prompt Example
4.3 Human Evaluation
We select a subset of size 50 (as done by Khondaker et al. (2023a) for Arabic) from the test dataset of each task. For human evaluation, two annotators (native Urdu speakers) are presented with anonymized outputs of Llama, Llama-FT, mT5 and GPT. They are asked to rank them from one to four based on the criteria in Appendix G, allowing multiple models to have the same rank. We prompt GPT, Llama and Claude with the same criteria, to assign a score to the outputs of the mentioned tasks, and then ranking is done based on this score. Since both GPT and Llama responses are being evaluated, we also include Claude as an additional evaluator because of the bias when LLMs assess their own outputs (Arif et al., 2024b). We compare ranking done by human annotators, and the three LLMs using Krippendorff’s alpha to determine the inter-rater reliability to determine the determine whether LLMs are effective evaluators for Urdu tasks.
5 Evaluation and Discussion
5.1 Classification
GPT Llama XLM-R Llama-FT
Task 0 3 6 CoT 0 3 6 CoT
Sentiment Analysis 90.98 91.17 94.90 94.60 87.88 60.62 59.71 57.54 92.90 95.30
Abuse Detection 86.27 89.01 88.71 87.62 44.73 65.85 71.64 75.82 90.92 89.15
Sarcasm Detection 58.17 69.18 66.56 65.47 44.94 34.25 50.67 49.75 84.37 81.48
Fake News Detection 78.88 75.95 78.14 76.45 66.36 63.48 71.45 40.98 84.99 72.14
Topic Classification 76.05 73.54 74.57 70.43 54.08 53.02 53.68 49.24 84.49 84.74
PoS Tagging 53.31 51.51 54.17 54.61 25.80 34.32 33.85 41.62 65.41 67.55
NER Tagging 61.96 62.18 62.98 63.95 42.13 40.80 45.29 53.96 70.41 90.90
Table 2: We report Macro-$F_{1}$ score for each classification task. GPT-4-Turbo and Llama-3-8b is evaluated in 0-shot, 3-shot, 6-shot and 6-shot with Chain-of-Thought settings. FT stands for fine-tuned.
We present the evaluation of the generalists (GPT, Llama) and the specialists (Llama-FT, XLM-R) for the classification tasks in Table 2. We observe that Llama-FT (fine-tuned) model achieves the highest scores in four tasks (sentiment analysis, topic classification, PoS tagging, and NER tagging), while XLM-R outperforms other models in three tasks (abuse detection, sarcasm detection, and fake news detection). GPT does not perform better than Llama-FT and XLM-R. For certain tasks like NER and PoS tagging, CoT reasoning leads to a better Macro-$F_{1}$ score. We now discuss the performance on each classification task in detail.
Sentiment Analysis. For sentiment analysis, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 95.30, outperforming XLM-R and GPT. GPT’s performance improves with more shots, reaching a Macro-$F_{1}$ of 94.90 with 6-shot learning. The CoT setting provides a slight increase to 94.60. However the Macro-$F_{1}$ for Llama decreases with few-shot and CoT prompting. The small difference in the Macro-$F_{1}$ between Llama-FT and GPT indicates the effectiveness of GPT for Urdu sentiment analysis without requiring task-specific fine-tuning.
Figure 8: Performance of GPT-CoT and Llama-FT on recognizing different entities in the NER task.
NER Tagging. For NER tagging, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 90.90, significantly outperforming XLM-R and GPT. The highest score achieved by GPT is 63.95, and by Llama is 53.96, both using CoT reasoning. We uncover the differences in accuracy for various entities for GPT and Llama-FT in Figure 8. We observe a drop in the recognition accuracy of Person entities by GPT, with 95 correctly recognized compared to 139 by Llama-FT. We notice a similar trend for Organization entities, with 70 classified correctly compared to 134 by Llama-FT. This suggests that NER fine-tuning on a specialized Urdu dataset improves the model’s ability to recognize Person and Organization entities in Urdu text.
Abuse Detection. In the task of abuse detection, XLM-R leads with a Macro-$F_{1}$ of 90.92, followed closely by Llama-FT at 89.15. GPT achieves its highest performance of 89.01 with 3-shot prompting, while Llama reaches a peak of 75.82 using CoT reasoning.
Sarcasm Detection. XLM-R performs best in sarcasm detection with a Macro-$F_{1}$ score of 84.37. Llama-FT also shows strong performance with 81.48. GPT achieves its best performance at 69.18 with 3-shot prompting, while Llama performs best at 50.67 with 6-shot prompting. The large difference between the Macro-$F_{1}$ scores of generalist and specialist models indicates that sarcasm detection in Urdu can be challenging without task-specific fine-tuning.
Fake News Detection. For fake news detection, XLM-R achieves the highest score with a Macro-$F_{1}$ of 84.99, while Llama-FT follows with 72.14. GPT’s performance is relatively consistent across different shot settings, with its highest score being 78.88 in the 0-shot setting. Llama’s highest score is 71.45 at 6-shot setting. The large margin by XLM-R indicates its effectiveness in discerning fake news.
Figure 9: Performance of GPT-CoT and Llama-FT on recognizing different parts of speech in the POS task.
GPT Llama mT5 Llama-FT
Task Metric 0 3 6 0 3 6
Question-Answering SQuAD-$F_{1}$ 66.28 72.55 72.78 69.89 71.65 71.62 69.66 70.42
Summarization ROUGE-L 22.54 23.35 23.34 25.76 25.84 26.17 30.72 31.01
Paraphrasing SacreBLEU 3.59 3.92 4.01 2.42 6.40 6.11 11.98 10.17
Transliteration SacreBLEU 30.93 32.38 32.09 12.37 18.51 21.35 40.23 37.95
Translation (en-ur) SacreBLEU 12.07 12.62 11.59 5.14 6.56 5.83 18.35 14.63
Translation (ur-en) SacreBLEU 16.29 18.05 19.18 7.83 13.11 12.32 21.55 29.18
Table 3: We report SQuAD-$F_{1}$, ROUGE-L or SacreBLEU depending on the generation tasks. GPT-4-Turbo and Llama-3-8b is evaluated in 0-shot, 3-shot and 6-shot setting.
Topic Classification. Llama-FT excels in topic classification with a Macro-$F_{1}$ score of 84.74, significantly higher than the other models. XLM-R also performs well with 84.49.
PoS Tagging. In PoS tagging, Llama-FT leads with a Macro-$F_{1}$ of 67.55, followed by XLM-R at 65.41. GPT’s highest score is 53.31 in the 0-shot setting and Llama’s highest score is 41.62 with CoT reasoning. To understand the significant difference in Macro-$F_{1}$ between GPT and Llama-FT, we study the individual class performance (Figure 9). We find that GPT struggles with correctly tagging proper nouns, pronouns, and auxiliaries, while Llama-FT is able to identify most of them correctly, suggesting that task-specific fine-tuning improves tagging performance for these parts of speech.
5.2 Generation
We present the evaluation of the generalists and the specialists for the generation tasks in Table 3 and Table 4. We observe that the fine-tuned Llama-FT model achieves the highest scores in summarization and translation from Urdu to English. On the other hand GPT shows a consistent performance across all the tasks with its best results often appearing in the 6-shot setting. The mT5 model also demonstrates strong performance in tasks such as transliteration and paraphrasing, benefiting from its extensive multilingual training on translation tasks. We now discuss the performance on each generation task in detail.
GPT’s Wins Llama’s Wins mT5’s Wins Llama-FT’s Wins
Task Annot. 1 Annot. 2 Annot. 1 Annot. 2 Annot. 1 Annot. 2 Annot. 1 Annot. 2
Summarization 41 34 9 17 7 12 4 11
Paraphrasing 42 39 12 15 4 3 4 4
Transliteration 40 38 5 6 13 21 19 21
Translation (en-ur) 46 46 3 1 7 10 13 10
Translation (ur-en) 40 45 12 10 14 15 18 15
AI Assistant 49 49 3 3 0 0 1 1
Table 4: We report the number of times each model is ranked 1 by each annotator for the generation tasks.
5.2.1 Quantitative Evaluation
Question-Answering. All the models show similar performance on question-answering task. GPT achieves the highest score with a SQuAD-F1 of 72.78 in the 6-shot setting. Llama closely follows with a score of 71.65 in 3-shot setting. Llama-FT has a lower score of 70.42, indicating that the translated UQA dataset is not very effective in improving the performance of the model. mT5 also performs well with a score of 69.66. The consistent improvement of GPT with increasing shots suggests its capability to use more context effectively.
Summarization. In the summarization task, Llama-FT achieves the highest ROUGE-L score of 31.01, outperforming Llama and GPT. mT5 also shows strong performance with a score of 30.72. Non-fine-tuned Llama outperforms GPT in all few-shot settings. GPT’s best performance is in the 3-shot setting with a score of 23.35 while Llama’s best is at 6-shot setting with a score of 26.17. The fine-tuning of Llama contributes to its superior performance in capturing and summarizing Urdu content effectively.
Paraphrasing. For paraphrasing, mT5 achieves the highest SacreBLEU score of 11.98. This indicates the advantage mT5 has due to its massive multilingual pre-training. Llama-FT follows with a SacreBLEU score of 10.17. Llama’s best score is 6.40 at 3-shot setting outperforming GPT’s best score of 4.01.
Transliteration. In the task of transliteration, mT5 leads with a SacreBLEU score of 40.23, followed by Llama-FT at 37.95. GPT’s performance peaks at 32.38 with 3-shot learning. Llama shows poor performace in this task with the highest score of 21.35 at 6-shot setting. Figure 10 shows the words with the highest mismatches in the transliterated text. To count these mismatches, we first tokenize the transliterated sentences and then count the instances where the predicted word differs from the ground truth. Smaller words such as “ye,” “ke,” “aik,” and “wo” pose challenges for GPT, resulting in higher mismatch counts. In contrast, mT5 demonstrates lower mismatches for these words.
Figure 10: Comparison between mismatches in the words predicted by mT5 and GPT for the transliteration task.
Translation (en-ur). For English to Urdu translation, mT5 achieves the highest SacreBLEU score of 18.35, indicating its proficiency in translating English to Urdu. Llama-FT follows with a score of 14.63. GPT’s performance is consistent, with its highest score being 12.62 in the 3-shot setting. Llama under-performs with a least highest score of 6.56 with 3-shot prompting. The superior performance of mT5 is likely due to the inclusion of high-quality Urdu data in the multilingual C4 corpus used for its pre-training Xue et al. (2021).
Translation (ur-en). In Urdu to English translation, Llama-FT excels with a SacreBLEU score of 29.18, outperforming other models. Surprisingly, contrary to the results in en-ur translation, mT5 shows a lower SacreBLEU of 21.55 compared to Llama-FT. GPT’s highest score is 19.18 in the 6-shot setting and Llama’s higest score is 13.11 in the 3-shot setting.
5.2.2 Qualitative Evaluation
AI Assistant. GPT outperforms all the models with 49 out of 50 wins as shown in Table 4. Llama-FT is ranked one 0 times according to both annotators and mT5 is ranked number one only 1 time. The low performance of fine-tuned models, such as Llama-FT and mT5, can be attributed to the fact that the fine-tuning dataset was translated from English.
Based on quantitative evaluation, GPT only performs the best in the question-answering task based on the quantitative evaluation provided in Table 3. However, Table 4 shows that GPT performs the best across all other tasks as well, according to qualitative evaluations by human annotators with more than 90% win-rate for most of the tasks. Figure 11 presents the number of times each rank was assigned to GPT for Urdu to English translation. Appendix F contains presents the rank counts for the other tasks.
Figure 11: Rank counts for GPT in human evaluation of Urdu to English translation.
Based on human evaluation LLama outperforms mT5 and even the fine-tuned Llama-FT on summarization, paraphrasing and AI assistant task. On transliteration and both translation tasks Llama-FT outperforms mT5. This discrepancy between quantitative and qualitative evaluations highlights an important issue: the translation of datasets from other languages may introduce inconsistencies or artifacts that impact the quantitative metrics, especially for tasks like summarization, paraphrasing, and story generation. This observation stresses the need for the creation of native Urdu datasets, rather than relying on translations of existing datasets. It also reflects the limitations and lower quality of the currently available Urdu datasets, which further exacerbates these challenges.
5.3 Factors Contributing to Improved Performance
Task-specific fine-tuning on Urdu datasets significantly improves the performance of LLMs on Urdu tasks, as evidenced by the superior results of fine-tuned Llama, XLM-R, and mT5 on both classification and generation tasks. Additionally, the inclusion of high-quality Urdu data in pre-training datasets, such as the clean multilingual C4 corpus, enhances downstream performance, with mT5 achieving higher SacreBLEU scores on transliteration, translation, and paraphrasing tasks compared to GPT and fine-tuned Llama. Providing more context in prompts also boosts performance, as shown by GPT’s improvement with increasing shots (0, 3, 6). Moreover, high-quality data for Indic languages like Hindi in pre-training improves performance on Urdu, especially for models like XLM-R, which excel in cross-lingual tasks, outperforming others in abuse detection, sarcasm detection, and fake news detection. The lack of native Urdu datasets poses a significant challenge for Urdu NLP, as most available datasets are translated from other languages.
5.4 Human Evaluators vs. LLMs
We compare human evaluation and LLM-based evaluation for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation, presented in Table 5.
Task A B C D
Summarization 0.684 0.504 0.502 0.624
Paraphrasing 0.710 0.592 0.471 0.565
Transliteration 0.694 0.510 0.253 0.603
Translation (en-ur) 0.728 0.592 0.392 0.515
Translation (ur-en) 0.730 0.474 0.307 0.510
AI Assistant 0.894 0.767 0.721 0.775
Table 5: A: Krippendorff’s alpha between annotator 1 and annotator 2. B: Alpha between both annotators and GPT. C: Alpha between both annotators and Llama. D: Alpha between both annotators and Claude.
For each task, the Krippendorff’s alpha value for human evaluation exceeds 0.67, which, according to Krippendorff’s interpretation is sufficient for a tentative conclusion to be drawn. Table 5 illustrates that the inclusion of GPT’s evaluation, Llama’s evaluation or Claude’s evaluation significantly reduces the alpha values meaning that the annotations done by GTP, Llama and Claude have a lower degree of agreement with the human annotators. For summarization, transliteration, Urdu-to-English translation, and AI assistant tasks, Claude has the highest agreement with humans, at 0.624, 0.604, 0.510, and 0.775, respectively. This showcases that the LLMs exhibit bias when the evaluated responses include their own generated content. For paraphrasing and Urdu-to-English translation tasks, GPT has the highest alpha values of 0.592 and 0.474, respectively. There is a higher agreement between the GPT rankings and human rankings as compared to the agreement between Llama rankings and human rankings.
6 Conclusion
In this paper, we present a comprehensive evaluation of generalist models and specialist models on 7 classification tasks and 7 generation tasks for Urdu NLP. Our evaluation covers prompting techniques such as few-shot, CoT reasoning as well as the fine-tuning of LLMs. We found that specialist models quantitatively outperformed generalist models on 12 out of the 14 tasks. The results highlight the importance of fine-tuning models to achieve higher performance in domain-specific applications in a low-resource setting. However, generalist models, such as GPT, showcased better performance in the human evaluation of the generation tasks, highlighting the importance of qualitative evaluation in accurately assessing model performance. We also performed a LLM based evaluation of the outputs of the models for the generation tasks. The low agreement between the rankings done by LLMs and human rankings shows that LLMs struggle when it comes to low-resource language understanding. Furthermore, the evaluation indicates a bias in LLMs when assessing responses that include their own outputs.
7 Future Work
One avenue for future research is to explore other strong general-purpose models (e.g., GPT-4o, GPT-4o1 and Claude) and expand the scope of the evaluation to more Urdu NLP tasks. Additionally, using Retrieval-Augmented Generation (RAG) to find examples from training dataset for few-shot prompting would be an interesting experiment to enhance the performance of generalist models. In conclusion, while specialist models currently hold an edge in classification tasks performance, generalist models’ adaptability and better performance in generation tasks remains valuable, and continuous advancements in LLMs promise further improvements for low-resource NLP.
8 Limitations
While our study provides valuable insights into the performance of generalist and specialist models for Urdu NLP tasks, it is important to acknowledge several limitations. The question-answering and sentiment analysis datasets used for training the specialist models are translated from English to Urdu. This translation process can introduce inaccuracies that may affect the models’ performance. Additionally, the lack of native, high-quality Urdu datasets poses a significant challenge. Translated datasets often fail to capture the linguistic and cultural nuances inherent to Urdu, which can impact both training and evaluation outcomes. The evaluation is conducted on a subset of 1000 data points for each task. While this size is manageable and allows for a cost-efficient evaluation, it may not be fully representative of the model’s overall performance. Although we use multiple annotators and calculate inter-rater reliability using Krippendorff’s alpha, there is still a degree of subjectivity that may influence the results.
9 Ethical Impact
In this paper we present a comprehensive evaluation of LLMs with the aim to enhance the accessibility of NLP applications for Urdu speakers. This has significant ethical implications, as it addresses the digital divide and promotes linguistic diversity in technology. Our findings indicate that specialist models perform better than the generalist models in most Urdu NLP tasks. Consequently, our work may inspire researchers to develop more resources for the Urdu language, including models and datasets.
The potential risks associated with the usage of LLMs include the amplification of existing biases present in their training data, which may lead to unfair and discriminatory outcomes (Ye et al., 2023). A comprehensive fairness evaluation of these models must be conducted before they are deployed for public use.
Acknowledgments
We are grateful for the time and effort put in by our research intern, Muhammad Suhaib Rashid who annotated our data and Mustafa Abbas who worked on creating Wiki-UQA dataset. We are also grateful to OpenAI for supporting our work through their Research Access Program.
Appendix A Implementation Details
A.1 Models
XLM-Roberta-large is availabe on Hugging Face [^9] https://huggingface.co/FacebookAI/xlm-roberta-large under MIT license. mT5-large is available on Hugging Face [^10] https://huggingface.co/google/mt5-large under Apache-2.0 license. Llama-3-8b is available on Hugging Face [^11] https://huggingface.co/meta-llama/Meta-Llama-3-8B under llama3 license. GPT-4 is available under proprietary licence. All models used in this paper comply with their respective license.
A.2 Datasets
Urdu IMDB sentiment analysis dataset OBDL license. Urdu NER dataset is availabe under MIT license. The abuse detection dataset by Akhter et al. (2020) and by Amjad et al. (2022), Sarcastic Tweets Dataset (Khan and Najeeb, 2023), fake news dataset by Amjad et al. (2020), Topic Classification dataset, and Universal Dependencies (Nivre et al., 2020) are available under CC BY 4.0.
UQuAD dataset is under CC0-1.0 while UQA (Arif et al., 2024a) is under CC BY 4.0. XSUMUrdu (Munaf et al., 2023) summarization dataset is also under CC BY 4.0 license. Paraphrasing dataset is under MIT license. Dakshina dataset (Roark et al., 2020) for transliteration is under CC BY-SA 4.0 and OPUS-100 (Zhang et al., 2020) (Tiedemann, 2012) for translation is under GPL-3.0 license. UrduAssistant (Khalil, 2023) dataset has a MIT license.
All datasets used in this paper comply with their respective license.
Appendix B Model Size and Budget
We fine-tuned XLM-Roberta-large for classification tasks which has 550 million parameters. We fine-tuned mT5-large for generation tasks which has 1.2 billion parameters. Llama-3-8b has 8 billion parameters and is fine-tuned for both generation and classification tasks.
Nvidia A100 80GB, Nvidia A100 40GB and Nvidia RTX 6000Ada 48GB were used for fine-tuning of the models. Infernece was done on Nvidia RTX 6000Ada 48GB and Nvidia RTX 4090 24GB. Total GPU time was approximately 200 hours.
Appendix C Human Annotators
There are two human annotators in this study: one is the author of this paper (Computer Science graduate), and the other is a research intern (Computer Science senior). Both annotators are native speakers of Urdu from Pakistan. The research intern was informed about how the data would be used for the evaluation of LLMs for Urdu.
Appendix D Dataset Size
This section provides details about the datasets used for the classification and generation tasks evaluated in this study. The test dataset sizes for each task are summarized in the tables below.
D.1 Classification
Table 6 shows the test dataset sizes for classification tasks including Sentiment Analysis, Abuse Detection, Sarcasm Detection, Fake News Detection, Topic Classification, Part-of-Speech (PoS) Tagging, and Named Entity Recognition (NER) Tagging.
Task Test Size
Sentiment Analysis 1000
Abuse Detection 567
Sarcasm Detection 1000
Fake News Detection 130
Topic Classification 1000
PoS Tagging 1000
NER Tagging 1000
Table 6: Test dataset size for each classification task
D.2 Generation
Table 7 shows the test dataset sizes for generation tasks including Summarization, Paraphrasing, Transliteration, English to Urdu Translation, and Urdu to English Translation.
Task Test Size
Question Answering 1000
Summarization 568
Paraphrasing 1000
Transliteration 1000
Translation (en-ur) 1000
Translation (ur-en) 1000
AI Assistant 50
Table 7: Test dataset size for each generation task
Appendix E Reproducibility and Hyperparameter
The table 8 presents the Lora configuration that we use for fine-tuning Llama.
Rank 16
Alpha 16
Target Modules q_proj, k_proj, v_proj, o_proj, down_proj, up_proj, gate_proj
Table 8: Lora config for Llama fine-tuning
The table 9 presents the training parameters, including the number of epochs and batch sizes, we use for fine-tuning XLM-R, mT5, and Llama.
Table 9: Training Parameters for Different Models. E represents number of epochs and B represents the batch size.
XLM-R mT5 Llama
Task E B E B E B
Sentiment Analysis 1 16 - - 4 1
Abuse Detection 4 16 - - 3 8
Sarcasm Detection 4 16 - - 3 8
Fake News Detection 6 8 - - 3 1
Topic Classification 4 32 - - 3 8
POS Tagging 7 16 - - 1 4
NER Tagging 10 16 - - 3 4
Question Answering - - 4 8 1 1
Summarization - - 1 4 2 1
Paraphrasing - - 3 3 4 4
Transliteration - - 5 8 2 4
Translation-en-ur - - 2 8 1 8
Translation - - 3 8 1 8
AI Assistant - - 2 8 2 8
Appendix F Human and LLM Evaluation
Figure 12 presents the rank distribution for all models across all generation tasks, as evaluated by two human evaluators: GPT-4-Turbo, Llama-3-8b, and Claude 3.5 Sonnet.
Appendix G Prompts and Evaluation Criteria
Table LABEL:fig:classprompts contains the CoT prompts used for the classification tasks. The prompts for classification without CoT are the same, except that the reasoning is not generated, and only the label is required in the output. Table LABEL:fig:genprompts contains the prompts used for the generation tasks. Table LABEL:fig:eval-criteria presents the criteria given to human evaluators, GPT and Llama for assessing the quality of outputs for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation. For each evaluation, the LLMs are asked to reason the answer to improve the scoring of the outputs.
Figure 12: Rank counts of models for all generation tasks.
Table 10: The table presents the CoT prompts used for classification tasks.
Classification Task CoT Prompt
Sentiment Analysis You are an Urdu sentiment classifier. The input text should be labeled according to its sentiment. The label list is: [’positive’, ’negative’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
Abuse Detection You are an Urdu abuse detector. The input text should be labeled according to whether it is abusive or not. The label list is: [’abusive’, ’not abusive’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
Sarcasm Detection You are an Urdu sarcasm detector. The input text should be labeled according to whether it is sarcastic or not. The label list is: [’sarcastic’, ’not sarcastic’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
Fake News Detection You are an Urdu fake news detector. The input text should be labeled according to whether it is fake news or not. The label list is: [’fake news’, ’not fake news’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
Topic Classification You are an Urdu topic classifier. The input text should be assign a label from the label list: [’business’, ’entertainment’, ’health’, ’politics’, ’science’, ’sports’, ’world’, ’other’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
PoS Tagging You are an Urdu part-of-speech tagger. The word wrapped in <hl> tag should be assigned a PoS tag from the label list: [’noun’, ’punctuation mark’, ’adposition’, ’number’, ’symbol’, ’subordinating conjunction’, ’adjective’, ’particle’, ’determiner’, ’coordinating conjunction’, ’proper noun’, ’pronoun’, ’other’, ’adverb’, ’interjection’, ’verb’, ’auxiliary verb’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
NER Tagging You are an Urdu named entity recognizer. The word wrapped in <hl> tag should be assigned a NER tag from the label list: [’time’, ’person’, ’organization’, ’number’, ’location’, ’designation’, ’date’, ’other’]
Use Chain-of-Thought (cot) to reason your answer.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"cot": ..., "label": ...}
Table 11: The table presents the prompts used for generation tasks.
Generation Task Prompt
Summarization You are a text summarizer. Your task is to summarize the given Pakistani Urdu text in 1 to 2 sentences. The summary should be in Urdu.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"summary": ...}
Paraphrasing You are a text paraphraser. Your task is to paraphrase the given Pakistani Urdu text. The paraphrased text should be in Urdu.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"paraphrase": ...}
Transliteration You are a machine transliterator. Your task is to transliterate the given Pakistani Urdu text to Roman Urdu.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"transliteration": ...}
Translation (en-ur) You are a machine translator. Your task is to translate the given English text to Pakistani Urdu.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"translation": ...}
Translation (ur-en) You are a machine translator. Your task is to translate the given Pakistani Urdu text to English.
ALWAYS RETURN JSON OBJECT IN FOLLOWING FORMAT ONLY:
{"translation": ...}
Table 12: The table presents the criteria used to evaluate the outputs of the generation task in the form of an LLM prompt.
Task Prompt/Criteria
Summarization You are a Pakistani Urdu language expert tasked with evaluating the quality of the summary produced by the summarization model. Score the given output with respect to the given input on a continuous score from 0 to 100 based on the following criteria:
1. Includes main points and key information from the original text
2. No grammatical errors
3. Conveys information in a brief manner
4. Gives correct information based on the original text
Think step by step and use reasoning. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}
Paraphrasing You are a Pakistani Urdu language expert tasked with evaluating the quality of paraphrased text produced by the paraphrasing model. Score the given output with respect to the given input on a continuous score from 0 to 100 based on the following criteria:
1. Retains the original meaning and key ideas
2. No grammatical errors
3. Use of different words and phrases than the original text
Think step by step and use reasoning. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}
Transliteration You are a Pakistani Urdu language expert tasked with evaluating the quality of transliterated text produced by the transliteration model. Score the given output with respect to the given input on a continuous score from 0 to 100 based on the following criteria:
1. Correctness of words (keeping in mind that different words can have different spelling)
2. Proper capitalization of words
Think step by step and use reasoning. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}
Translation (en-ur) You are a Pakistani Urdu language expert tasked with evaluating the quality of translated text produced by the translation model. Score the given output with respect to the given input on a continuous score from 0 to 100 based on the following criteria:
1. Conveys the meaning of the original text without omissions or additions
2. No grammatical errors
3. Retains the style and tone of the original text
Think step by step and use reasoning. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}
Translation (ur-en) You are a Pakistani Urdu language expert tasked with evaluating the quality of translated text produced by the translation model. Score the given output with respect to the given input on a continuous score from 0 to 100 based on the following criteria:
1. Conveys the meaning of the original text without omissions or additions
2. No grammatical errors
3. Retains the style and tone of the original text
Think step by step and use reasoning. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}
AI Assistant Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, please rate the response on a scale of 0 to 100. ALWAYS RETURN JSON OBJECT IN THE FOLLOWING FORMAT ONLY:
{"reasoning": ..., "score": ...}